OpenStudy (anonymous):

a person 2 m tall casts a shadow 5 m long. at the same time, a building casts a shadow 24 m long. how tall is the building. help me please

OpenStudy (anonymous):

Set up a proportion: 2m (tall) / 5m (shadow) = x (tall) / 24m (shadow) 5x = 2(24) {cross-multiplied} 5x = 48 {multiplied 2 by 24 x = 48/5 {divided both sides by 5} x = 9.6 m tall {the building}